### v3

# R2C2-Coder: Enhancing and Benchmarking Real-world Repository-level Code Completion Abilities of Code Large Language Models

Abstract
Recently, repository-level code completion has drawn more attention in modern software development, and several baseline methods and benchmarks have been proposed. However, existing repository-level code completion methods often fall short of fully using the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies. Besides, the existing benchmarks usually focus on limited code completion scenarios, which cannot reflect the repository-level code completion abilities well of existing methods. To address these limitations, we propose the R2C2-Coder to enhance and benchmark the real-world repository-level code completion abilities of code Large Language Models, where the R2C2-Coder includes a code prompt construction method R2C2-Enhance and a well-designed benchmark R2C2-Bench. Specifically, first, in R2C2-Enhance, we first construct the candidate retrieval pool and then assemble the completion prompt by retrieving from the retrieval pool for each completion cursor position. Second, based on R2C2-Enhance, we can construct a more challenging and diverse R2C2-Bench with training, validation and test splits, where a context perturbation strategy is proposed to simulate the real-world repository-level code completion well. Extensive results on multiple benchmarks demonstrate the effectiveness of our R2C2-Coder.
footnotetext: * First three authors contributed equally.footnotetext: Corresponding Author: Jiaheng Liu.
1 Introduction
Large Language Models for Code (Code LLMs), such as Codex (Chen et al., 2021), CodeGen (Nijkamp et al., 2023a, b), and StarCoder (Li et al., 2023), have shown their power to enhance developer productivity as their promising results in code completion (Athiwaratkun et al., 2023; Austin et al., 2021; Chen et al., 2021; Lu et al., 2021). Besides, to evaluate these models in real-world scenarios, multiple code completion benchmarks are proposed. Recently, the challenging repository-level code completion has drawn more attention, where the extensive cross-file dependencies (i.e., contextual information from other code files within the same repository) are used to enhance the completion performance. Therefore, a direct but complex solution is to enhance the context length of LLMs (Chen et al., 2023a, b; Peng et al., 2023). Nonetheless, increasing the context length significantly raises costs and is not affordable, especially concerning the large number of files in a given repository. Hence, several recent baseline methods (Ding et al., 2023; Liu et al., 2023b) apply the widely-used Retrieval-Augmented Generation (RAG) strategy by retrieving the most relevant code snippets to improve the repository-level code completion.
However, this vanilla strategy exhibits significant limitations. First, it usually preprocesses the source code files into code snippets under a repository in Fig. 1, and performs retrieval on these code fragments for retrieval query. However, these fragments only preserve local and discrete information and fail to cover more diverse and fundamental contexts of program semantics, such as the intricate network of dependencies, shared utility functions, definition of the input and output parameters, etc. Besides, as retrieved code snippets based on retrievers (e.g., BM25 Robertson et al. (2009)) usually contain irrelevant or noisy contexts, which may affect the completion results, it is important to evaluate and improve the robustness of the noisy contexts for Code LLMs. Moreover, we observe that the existing repository-level code completion benchmarks focus on limited completion scenes and cannot fully simulate real-world usage well (See Section 4).
Figure 1: Examples of code snippets.
To address these limitations, we propose the R2C2-Coder to enhance and benchmark the Real-world Repository-level Code Completion abilities of Code LLMs, where a code prompt construction method (i.e., R2C2-Enhance) and a challenging benchmark (i.e., R2C2-Bench) are proposed.
First, R2C2-Enhance includes two stages (i.e., candidate retrieval pool construction, and completion prompt construction). In the retrieval pool construction stage, we first build the retrieval pool using the abstraction context based on the parser generator tool (e.g., Tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/) and the snippet context, where the abstraction context aims to represent the coarse-grained global information of each programming file and the snippet context is built by extracted code fragments to provide the fine-grained local information for completion. Then, in completion prompt construction stage, for the current cursor position [^2] The cursor position denotes the location, where code completion is about to be triggered., we first build the retrieval query and produce the retrieved contexts by performing the retrieval between the retrieval query and the retrieval pool, where the retrieved contexts are then combined with the current code file to generate the completion prompt. Finally, the completion prompt is sent to Code LLMs to generate the completion response.
Second, based on R2C2-Enhance, we can easily build a real-world repository-level code completion benchmark called R2C2-Bench, which includes the training, validation and test splits in 4 languages. Note that when producing R2C2-Bench, we also apply a context perturbation strategy to simulate more diverse and challenging code completion samples for generating the completion prompt.
By performing experiments on the validation and test splits of R2C2-Bench, we observe that our R2C2-Enhance achieves significant improvements when compared with the existing methods without training. Besides, when we further fine-tune the Code LLMs using the training split of R2C2-Bench, better results are obtained on the validation and testing splits of R2C2-Bench. Moreover, we also validate the generalization abilities of R2C2-Enhance on the widely-used CrossCodeEval benchmark.
The contributions of our R2C2-Coder are as follows:
  • We investigate the limitations of existing repository-level code completion (e.g., lack of sufficient context and benchmark with limited scenarios) and propose the R2C2-Coder including R2C2-Enhance and R2C2-Bench to enhance and benchmark real-world repository-level code completion abilities of Code LLMs.
  • For R2C2-Enhance, we propose to construct the candidate retrieval pool with abstract and snippet contexts and generate the completion prompt using context retrieval and prompt assemble. Based on R2C2-Enhance, we build a new repository-level code completion benchmark called R2C2-Bench with training, validation, and testing splits, where a context perturbation strategy is used to simulate the real-world completion scenes better.
  • Comprehensive experimental results on multiple benchmark datasets demonstrate the effectiveness and efficiency of our R2C2-Coder.
2 Related Works
Code Large Language Models. The successive achievements of generative language modeling (Black et al., 2022, 2021; Brown et al., 2020; Radford et al., 2019; Touvron et al., 2023; Wang and Komatsuzaki, 2021; Bai et al., 2024; Wu et al., 2024; Liu et al., 2024; Sun et al., 2024; Guo et al., 2024b; Wang et al., 2023b; Guo et al., 2023; Zhang et al., 2024; Du et al., 2024) have inspired extensive studies on generative AI for software engineering. While closed-source models (Achiam et al., 2023; Chen et al., 2021; Chowdhery et al., 2023) achieve significant dominance in the benchmark leaderboards (Chen et al., 2021; Hendrycks et al., 2020; Liu et al., 2023a), their inaccessible model checkpoints and source code are detrimental to subsequent innovation. In contrast, plenty of open-source models, for instance, Code T5 (Wang et al., 2023a, 2021), CodeGen (Nijkamp et al., 2023a, b), StarCoder (Allal et al., 2023; Li et al., 2023; Lozhkov et al., 2024), Code Llama (Roziere et al., 2023) and DeepSeekCoder (Guo et al., 2024a), have been proposed and have substantially driven the development of code intelligence.
Repository-level Code Completion. Existing methods are usually based on a retrieval augmented generation (RAG), and we categorize related works into ranking-based and fusion-based methods. Ranking-based approaches (Liao et al., 2023; Pei et al., 2023; Phan et al., 2024; Zhang et al., 2023) explicitly select those code snippets in the repository that are highly similar to the incomplete code. In contrast, fusion-based methods (Ding et al., 2022; Shrivastava et al., 2023a, b) focus on organizing the repository-level context to be jointly modeled with the language model and allow the model to automatically select the most relevant information.
Code Benchmark. Program synthesis is to prompt code models to solve programming problems using the input description. The generated code snippet should be functional to pass all test cases (Athiwaratkun et al., 2023; Austin et al., 2021; Chen et al., 2021; Gu et al., 2024; Lai et al., 2023; Liu et al., 2023a; Yu et al., 2024). Besides, numerous benchmarks can be used to comprehensively evaluate code models, such as code translation (Jiao et al., 2023; Yan et al., 2023; Zhu et al., 2022), code retrieval (Huang et al., 2021; Husain et al., 2019; Li et al., 2024; Lu et al., 2021), and vulnerability repair (Huq et al., 2022; Prenner and Robbes, 2023; Richter and Wehrheim, 2022; Tian et al., 2024). Recently, several benchmarks have been proposed on repository-level code completion (Agrawal et al., 2023; Allal et al., 2023; Bairi et al., 2023; Ding et al., 2022; Liu et al., 2023b; Pei et al., 2023; Shrivastava et al., 2023b, a; Zhang et al., 2023). For example, RepoBench (Liu et al., 2023b) measures the retrieval and completion performance of the system. CrossCodeEval (Ding et al., 2023) evaluates three cross-file prompting levels: in-file, retrieval, and retrieval within references. However, RepoBench only supports two programming languages and neglects the FIM setting. CrossCodeEval only considers explicit file imports, resulting in few cross-file dependencies and low multi-line prediction difficulty.
3 R2C2-Coder
In this section, we discuss the details of R2C2-Coder, which includes a code prompt construction method R2C2-Enhance and a new benchmark R2C2-Bench. For R2C2-Enhance, in Fig. 2, we first build the candidate retrieval pool based on the abstract context and the snippet context to produce the coarse-grained global information of each programming file and the fine-grained local information of each code fragment, respectively. Then, for the current triggered cursor location, we produce the retrieval query and obtain the retrieved contexts based on the completion query and the candidate context pool, where the retrieved contexts and the current in-file context are used to construct the completion prompt. Based on R2C2-Enhance, we can easily produce the R2C2-Bench, where a context perturbation strategy is used to generate more diverse and challenging completion scenes.
Figure 2: Overview of our R2C2-Enhance. For the current completion cursor position, we first generate the retrieval query using the prefix and suffix contexts. Then, we perform context retrieval between the retrieval query and the pre-constructed candidate retrieval pool to produce the retrieved contexts. After that, we use the in-file context of the current code and the retrieved contexts to assemble the completion prompt, which is then sent to LLMs to generate the completion response.
3.1 R2C2-Enhance
3.1.1 Construction of the Candidate Retrieval Pool
Given a repository, we extract abstract and snippet contexts to construct the candidate retrieval pool.
Figure 3: An example of the abstract syntax tree generated by the Tree-sitter tool.
Abstract Context Extraction. In Fig. 2, we propose to use the parser generator tool (i.e., Tree-sitter [^3] https://tree-sitter.github.io/tree-sitter/) to extract the abstract context for each file in the repository. Specifically, we use Tree-sitter to generate the abstract syntax tree in Fig. 3, and then preserve these nodes to represent the declarations (e.g., functions, variables, classes etc.). After that, for each node, we remove the redundant information (e.g., middle variables, empty lines, comments etc.) and only preserve the information related to the definition for each declaration. Note that the extraction process is easy and efficient (See Appendix B.1 for more details.). Moreover, in Fig. 9 of Appendix B.1, we provide examples of the extracted abstract contexts and observe that the abstract context mainly provides the coarse-grained global information for each code file.
Snippet Context Extraction. For the snippet context, we iteratively scan the files in the repository and extract contiguous $M$ lines (in all our experiments, we set $M=10$ follow Ding et al. (2023)) of overlapping code fragments, which are the candidates for context retrieval.
3.1.2 Construction of the Completion Prompt
Context Retrieval. In Fig. 2, for the current cursor position, we first build the retrieval query based on the prefix and suffix contexts. Specifically, for the current cursor position, we first use the previous $P$ lines and subsequent $S$ lines as the prefix and suffix contexts, respectively. Therefore, the retrieval query usually has $P+S$ lines. Note that if the prefix and suffix have one overlapped line, the retrieval query has P+S1P+S-1 lines. Then, we use the retriever (e.g., BM 25 (Robertson et al., 2009)) to calculate the similarity between the retrieval query and the candidate retrieval pool. After that, we can select the retrieved contexts based on the similarities. Note that in the retrieval process, we remove the abstract context and snippet context of the current code file, where the cursor position is located.
Prompt Assemble. We assemble the completion prompt by using the current code file, and retrieved contexts. Specifically, we first follow Ding et al. (2023) to maintain all context of the current code file as the in-file context, where the completion cursor position is located. Then, we preserve the top-$K$ similar abstract contexts as the coarse-grained global structure information, After that, we append the most relevant snippet context based on the similarity scores until the maximum number of tokens is $N$, where we set $N$ as 4,096 by default. Finally, we can obtain the assembled prompt with $N$ tokens, which will be sent to the Code LLMs for generating the completion results.
3.2 R2C2-Bench
In this section, we provide the details of our R2C2-Bench with training, validation and test splits.
3.2.1 Dataset Generation
Initial Repository Collection. We collect permissively licensed repositories from GitHub. Specifically, following (Ding et al., 2023), we contain four languages and keep only repositories with the number of stars >=3>=3. Meanwhile, we exclude repositories that contain fewer than 10 or more than 50 source code files, and we eliminate any repositories that have at least one source code file identical to any file found in the Stack dataset (Kocetkov et al., 2022). As a result, we ended up with 54972, 51796, 49790, and 35410 repositories for Python, Java, TypeScript, and C#, respectively.
Completion Cursor Position Generation. Based on R2C2-Enhance, given completion cursor position, we can easily generate the completion prompt. Therefore, it is critical to generate informative and sufficient cursor positions for our R2C2-Bench. Specifically, in Fig. 3, for each file in the repository, we reuse the abstract syntax tree generated in the abstract context extraction stage, and randomly choose one node of the tree as the cursor position to be triggered for completion. Then, for the selected node, we can easily generate one completion prompt sample for our R2C2-Bench.
Feature Python Java TypeScript C#
# Repositories 260 397 451 245
# Files 2500 2500 2500 2500
# Examples 4394 6900 6506 5028
Avg. # cross files in prompt 6.4 7.8 6.8 10.1
Avg. # lines in prompt 320.5 333.7 365.7 373.8
Avg. # tokens in prompt 3192.25 3262.3 3200.52 3313.23
Avg. # lines in reference 1.73 1.75 2.02 1.78
Avg. # tokens in reference 17.47 14.22 15.00 13.32
Table 1: The statistics of the testing split of R2C2-Bench.
Context Perturbation. In real-world repository-level code completion scenes, we cannot always retrieve the relevant context well. For example, in certain software development scenes, we create code files to implement different usages under a repository. However, the relationships between these code files are relatively low, which indicates that the constructed completion prompt contains many irrelevant contexts. Therefore, to simulate the real-world repository-level code completion scene and improve the semantic variety of the completion prompt, we introduce a so-called context perturbation strategy. Specifically, for $Q\%$ completion cursor position (we set $Q=10$ by default), we first sort the contexts based on the similarities and follow the uniform distribution to determine the proportion $R\%$. Then we randomly discard top $R\%$ contexts with high similarities. After that, we perform the context retrieval based on the retrieval query and the remaining contexts, and the generated retrieved contexts are then combined with the current file to produce the completion prompt as discussed in Sec. 3.1.2.
Based on the above meticulous procedure, we can produce sufficient completion prompt samples using these repositories, and randomly select 400,000 as training samples, where each language has 100,000 samples. For the validation and testing samples, to ensure the quality of the dataset, we apply a series of rule-based and model-based post-processing filter strategies discussed in Appendix B.2.
3.2.2 Dataset Statistics
The statistics of R2C2-Bench are shown in Table 1444We use the StarCoder tokenizer (Li et al., 2023) to compute the number of tokens.. We have provided the statistics of CrossCodeEval and validation split of R2C2-Bench in Appendix B.4
Model Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
StarCoder-7B 17.0 49.5 21.6 55.9 18.1 53.6 17.6 51.0 18.6 52.5
+ R2C2-Enhance 21.4 46.2 27.6 51.9 24.8 52.8 25.2 45.8 24.8 49.2
+ R2C2-Enhance w.o. Abs. 21.5 46.2 26.1 51.4 24.6 52.8 24.9 45.7 24.3 49.0
+ R2C2-Enhanced Tuning 34.0 65.0 45.3 74.4 37.0 68.9 45.0 71.6 40.3 70.0
Code Llama-7B 13.7 44.1 20.6 54.5 17.4 51.2 16.0 48.4 16.9 49.5
+ R2C2-Enhance 16.3 39.0 22.3 45.0 22.1 46.5 18.8 36.0 19.9 41.6
+ R2C2-Enhance w.o. Abs. 16.2 38.9 21.6 44.8 22.0 46.4 19.1 36.9 19.7 41.7
+ R2C2-Enhanced Tuning 34.5 66.5 47.4 77.2 40.2 71.7 48.0 74.7 42.5 72.5
DeepSeekCoder-6.7B 19.4 52.5 24.2 59.3 22.0 58.8 20.7 54.1 21.6 56.2
+ R2C2-Enhance 25.8 51.3 31.8 56.5 29.7 57.8 33.1 53.6 30.1 54.8
+ R2C2-Enhance w.o. Abs. 25.8 51.3 30.7 55.6 29.6 58.0 31.9 52.5 29.5 54.3
+ R2C2-Enhanced Tuning 38.0 69.6 49.2 78.0 43.0 73.8 52.3 77.5 45.6 74.7
Table 2: Exact match (%) and edit similarity (%) performance on R2C2-Bench. The baseline results are produced under the “In-file Context Only” setting while “+R2C2-Enhance” means the input is prepended with retrieval results fetched by R2C2-Enhance at inference. In “+ R2C2-Enhance w.o. Abs.”, we report results without using abstract contexts in the candidate retrieval pool. “R2C2-Enhance w/ F.T.” means we fine-tune these LLMs on the training set of R2C2-Bench with R2C2-Enhance.
4 Experiments
We conduct experiments on CrossCodeEval (Ding et al., 2023), CrossCodeEval+, and R2C2-Bench with three popular Code LLMs (i.e., StarCoder-7B (Li et al., 2023), DeepSeekCoder-7B (Guo et al., 2024a) and Code Llama-7B (Roziere et al., 2023)) (See Appendix C.1 for more details). For CrossCodeEval+, we collect the original repositories of CrossCodeEval (Ding et al., 2023) then parse and extract the cross-file dependencies using R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in the original CrossCodeEval (See Appendix C.2 for more details).
4.1 Experimental Setup
In-file Context Only. We only provide the original code file, where the cursor position is located. Without any explicit cross-file context as input, the model should rely on its knowledge-based reasoning capabilities to accomplish the code generation process.
Retrieval in CrossCodeEval. This setup is similar to the “+ Retrieval” setting of CrossCodeEval (Ding et al., 2023), where the retrieval candidates are constructed by exhaustively scanning files in the same repository and extracting contiguous-$M$-line code fragments ($M=10$). Then, the top 5 similar candidates and the test query assemble the prompt. Note that for better baseline performance, we do not truncate cross-file contents into a total length of 512 tokens as (Ding et al., 2023).
Retrieval with R2C2-Enhance. Based on R2C2-Enhance in Sec. 3.1, given a repository, we first construct the candidate retrieval pool via abstract and snippet context extraction. Then, we build the retrieval query including prefix length $P=5$ and suffix length $S=5$. After that, candidates are concatenated to the beginning of the in-file context based on their Jaccard similarity (Jaccard, 1912) in descending order, until the cumulative length, including the in-file context, reaches the maximum token limit of $N=4096$. Note that we set the number of the abstract context $K$ as 3.
R2C2-Enhanced Tuning. To further boost the performance, we fine-tune code LLMs on the training split of R2C2-Bench including 400,000 files mentioned in Sec. 3.2.1.
Evaluation Metrics. Following CrossCodeEval (Ding et al., 2023), we evaluate code LLMs with code match metrics including exact match (EM) and edit similarity (ES). Exact match expects that the prediction and the real code are extremely identical, while edit similarity flexibly allows the generated samples to match the original code through a certain number of editing operations.
4.2 Results
We present results on R2C2-Bench in Table 2, CrossCodeEval (Ding et al., 2023) in Table 3, and CrossCodeEval+ in Table 4. Based on Table 3 and Table 4, cross-file context is undoubtedly efficacious, while retrieval with R2C2-Enhance surely exhibits its superiority compared with retrieval method in CrossCodeEval. Moreover, fine-tuning strategy significantly improves the performance on both benchmarks. Besides, when abstract contexts are removed from the candidate retrieval pool, these code LLMs suffer from degraded performance. Notably, in Table 2, after applying R2C2-Enhanced tuning, the Code Llama-7B, which is the worst in the in-file setting, outperforms the StarCoder-7B. Overall, we could say that there exists a performance promotion path, “In-file” $\rightarrow$ “+ R2C2-Enhance w.o. Abs.” $\rightarrow$ “+ R2C2-Enhance” $\rightarrow$ “+ R2C2-Enhanced Tuning”.
Model Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
StarCoder-7B 24.3 51.3 27.7 61.4 30.8 67.1 44.7 69.8 30.9 62.1
+ Retrieval 27.0 51.0 33.6 62.9 33.1 66.7 47.1 67.7 34.1 61.9
+ R2C2-Enhanced Tuning 34.2 64.3 40.2 73.7 37.8 72.4 50.6 77.8 39.6 71.5
Code Llama-7B 22.0 46.8 28.6 60.4 10.1 63.0 44.5 71.3 23.4 59.6
+ Retrieval 23.0 45.3 33.8 61.3 10.7 63.1 49.3 70.1 25.8 59.2
+ R2C2-Enhanced Tuning 34.7 65.4 41.0 72.7 11.4 69.5 53.2 77.8 31.5 70.6
DeepSeekCoder-6.7B 25.9 51.8 31.3 62.5 33.8 67.5 42.9 67.6 32.8 62.2
+ Retrieval 28.0 52.6 36.2 62.3 35.6 66.7 48.4 67.7 35.9 62.1
+ R2C2-Enhanced Tuning 37.7 67.3 45.1 75.3 42.0 75.7 59.3 81.8 44.6 74.5
Table 3: Exact match (%) and edit similarity (%) performance on CrossCodeEval.
Model Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
StarCoder-7B 24.3 51.3 27.7 61.4 30.8 67.1 44.7 69.8 30.9 62.1
+ R2C2-Enhance 30.9 51.9 38.1 63.6 35.5 67.7 68.4 78.2 40.7 64.5
+ R2C2-Enhance w.o. Abs. 30.9 51.8 35.3 62.6 34.5 67.6 65.2 77.1 39.2 64.0
+ R2C2-Enhanced Tuning 41.8 68.5 46.2 76.0 40.7 74.4 69.3 86.5 47.3 75.3
Code Llama-7B 22.0 46.8 28.6 60.4 10.1 63.0 44.5 71.3 23.4 59.6
+ R2C2-Enhance 23.6 42.9 35.6 58.5 10.6 61.6 67.8 78.5 29.7 58.9
+ R2C2-Enhance w.o. Abs. 23.9 43.0 34.1 58.9 11.0 62.3 65.4 78.0 29.1 59.2
+ R2C2-Enhanced Tuning 43.0 71.8 48.8 76.8 13.0 73.2 70.3 87.0 39.0 76.1
DeepSeekCoder-6.7B 25.9 51.8 31.3 62.5 33.8 67.5 42.9 67.6 32.8 62.2
+ R2C2-Enhance 32.7 54.0 41.6 64.6 38.8 67.9 69.0 78.8 43.1 65.4
+ R2C2-Enhance w.o. Abs. 32.5 53.7 38.8 63.5 37.8 67.5 65.7 77.5 41.5 64.7
+ R2C2-Enhanced Tuning 46.9 73.5 53.2 79.3 45.8 78.2 78.9 89.9 53.6 79.3
Table 4: Exact match (%) and edit similarity (%) performance on CrossCodeEval+.
Ablation Models $P$ $S$ Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
StarCoder-7B + R2C2-Enhanced Tuning 5 5 40.8 68.5 42.0 74.1 35.6 68.1 50.0 76.6 42.1 71.8
    Prefix & Suffix 5 1 38.0 64.0 41.6 71.8 31.6 64.7 48.0 74.2 39.8 68.7
5 3 40.0 66.8 41.6 73.8 32.0 65.5 49.2 75.6 40.7 70.4
    Prefix Only 10 - 39.2 68.7 39.2 71.0 32.4 67.5 47.6 74.5 39.6 70.4
    w.o. Snippet Context 1 1 24.8 59.1 36.4 71.5 26.8 64.4 36.8 70.3 31.2 66.3
3 3 26.4 59.7 36.0 71.5 25.6 64.0 35.6 70.0 30.9 66.3
5 5 25.6 59.0 37.6 71.4 24.8 64.0 38.4 70.8 31.6 66.3
10 10 24.0 58.1 36.4 71.5 25.2 64.2 35.2 68.4 30.2 65.5
    w.o. Abstract Context 1 1 38.0 65.9 37.6 69.3 31.6 65.3 42.8 71.4 37.5 68.0
3 3 37.2 62.9 39.2 68.3 31.6 62.5 42.4 69.3 37.6 65.7
5 5 37.2 65.1 38.8 69.8 33.2 65.2 45.2 73.3 38.6 68.4
10 10 37.6 66.1 37.2 68.2 32.4 64.8 44.8 72.8 38.0 68.0
Table 5: Exact match (%) and edit similarity (%) performance on R2C2-Bench for StarCoder-7B. During retrieval, $P$ prefix lines and $S$ suffix lines are concatenated as the query. In “Prefix Only”, no suffix lines are taken as the query. “w.o.” stands for removing the snippet contexts or abstract contexts at the fine-tuning and the inference stages simultaneously.
4.3 Analysis
Ablations. We conduct ablation studies on R2C2-Enhance and present the results in Table 5. We fine-tune StarCoder-7B (Li et al., 2023) with R2C2-Enhance and report the performance on the validation set of R2C2-Bench. Meanwhile, we conduct experiments on the retrieval manner, where we replace the FIM mode with “prefix-only”. We could find that suffix contexts are necessary as the results of $P=5,S=3$ even surpass those of $P=10$. Additionally, we remove either snippet contexts (“w.o. Snippet Contexts”) or abstract contexts (“w.o. Abstract Contexts”) simultaneously in fine-tuning and testing. Results on various prefix lengths ($P$) and suffix lengths ($S$) are further reported for the ablation model. Despite the larger impact without using snippet contexts, abstract contexts still deliver 3.5% and 3.4% performance gains. Furthermore, a case study in Fig. 6 shows how the code LLMs benefit from abstract context extraction.
Context Perturbation Helps Cope with Confusing Contexts. In real-world scenarios, retrievers would produce low-quality results as shown in Fig. 7. To enhance the model robustness on irrelevant contexts, we select $Q\%$ of the cursors and randomly remove the top similar contexts during R2C2-Enhanced tuning. Fig. 4 visualizes how the performance of StarCoder-7B on EM and ES scales when using different perturbation rates for R2C2-Enhanced tuning. The perturbation rate $Q\%$ for fine-tuning varies in $0\%,10\%,20\%,50\%,$ and $80\%$, while the perturbation rate for the validation set is always kept as 10%. The best results are achieved under perturbation rate $Q\%=10\%$, which is also the default setting of R2C2-Enhanced tuning for Table 2 , 3 and 4.
Similarity Metric Python Java TypeScript C# Average
Execution Time
EM ES EM ES EM ES EM ES EM ES
 Jaccard 40.8 68.5 42.0 74.1 35.6 68.1 50.0 76.6 42.1 71.8 2.941
 BM25 38.4 64.6 40.8 71.5 31.6 65.6 50.0 75.4 40.2 69.3 3.142
 UniXCoder 39.2 66.0 40.8 70.9 33.6 66.1 49.2 75.9 40.7 69.7 250.9
Table 6: Performance of StarCoder-7B with different similarity metrics on the validation set of R2C2-Bench. The retrieval time consumption (ms) per sample is further reported.
Figure 4: Performance of StarCoder-7B + R2C2-Enhanced Tuning at various perturbation rates on the validation set of R2C2-Bench.
Lexical Retrievers are Cost-effective. For ranking-based code retrieval, several approaches could be adopted as the similarity metric. We categorize the metrics into lexical ones, including Jaccard similarity (Jaccard, 1912) and BM25 (Robertson et al., 2009), and neural ones such as UniXcoder (Guo et al., 2022). In real-world scenarios, we often need to make a trade-off, as lexical retrievers could save more response time (RT) while neural retrievers would theoretically achieve higher performance. To find the most suitable retriever for R2C2-Enhance, we evaluate the code match performance and retrieval duration with Jaccard similarity, BM25, and UniXCoder. Table 6 presents the code match performance and retrieval duration of each retriever. Although UniXCoder outperforms BM25 on code match, its time consumption is extremely worse than lexical retrievers. We find that Jaccard similarity is the best retriever regarding both effectiveness and efficiency. Thus, we choose Jaccard similarity by default. See Appendix C.4 for details of execution time benchmarking.
Scalability of Model Performance. We present how the performance of StarCoder scales w.r.t model parameters in Table 7. We find that almost no performance gain is obtained for StarCoder-3B when using R2C2-Enhance with abstract and snippet contexts. After R2C2-Enhanced tuning, StarCoder-3B can outperform the inference-only results of StarCoder-7B. In Table 8 we further evaluate the fine-tuned StarCoder-7B when using different sizes of tuning datasets based on our R2C2-Enhanced tuning. We observe that when increasing from 10K samples to 100K samples per language, better results are obtained. When continually increasing, the results are relatively stable. Thus, we choose 10K samples per language as the default training set size of R2C2-Bench.
Scalability on Multi-line Predictions. In Fig. 5, we provide a detailed analysis of the scalability on multi-line predictions. Specifically, in Fig. 5(a), we provide the counts of completion references with different lines and observe that our R2C2-Bench has more multi-line completion scenes. Then, in Fig. 5(b) and Fig. 5(c), we provide the results of StarCoder-7B with R2C2-Enhanced Tuning on CrossCodeEval+ and R2C2-Bench, and observe that the scores on R2C2-Bench are lower than CrossCodeEval+ a lot, specifically in multi-line scenes, which means that our R2C2-Bench is more challenging than the CrossCodeEval+.
Limitations of CrossCodeEval. We summarize three major drawbacks of CrossCodeEval (Ding et al., 2023) as follows: 1) The retrieval results provided in CrossCodeEval are produced by sliding-window-based searching, which could potentially miss long-term dependencies that exceed the length of $M$. In contrast, abstract extraction expands the reception field of code LLMs with coarse-grained global information. 2) CrossCodeEval only considers explicit dependencies (i.e., file imports), but there could be implicit references during software development. For example, in Java, usages in the same package do not require the import statement. As shown in Fig. 8, such implicit dependencies could not be detected by Ding et al. (2023), but are successfully covered by our R2C2-Enhance during constructing CrossCodeEval+. 3) Compared with CrossCodeEval, our R2C2-Bench provides more complex cross-file dependencies (see “Avg. # cross files in prompt” in Table 1 and Table B.4) and more difficult multi-line test samples in Fig. 5).
(a) Multi-line Reference Counts
(b) Exact Match
(c) Edit Similarity
Figure 5: (a) Statistics of references from 1 to 5 lines in R2C2-Bench and CrossCodeEval+. (b) Exact Match of StarCoder-7B w/ R2C2-Enhanced tuning on R2C2-Bench and CrossCodeEval+ when expected output varies from 1 to 5 lines. (c) Edit Similarity of StarCoder-7B w/ R2C2-Enhanced tuning on R2C2-Bench and CrossCodeEval+ when expected output varies from 1 to 5 lines.
Model Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
StarCoder-3B 10.8 38.6 18.0 51.1 12.4 49.0 14.8 50.0 14.0 47.2
+ R2C2-Enhance 16.8 34.1 18.0 39.7 16.0 39.0 12.8 31.2 15.9 36.0
+ R2C2-Enhanced Tuning 40.0 64.4 37.2 68.0 28.0 61.8 45.6 71.0 37.7 66.3
StarCoder-7B 15.2 45.2 26.0 59.1 14.4 52.4 16.0 52.6 17.9 52.3
+ R2C2-Enhance 23.2 46.4 30.8 56.8 25.2 52.6 26.4 48.9 26.4 51.2
+ R2C2-Enhanced Tuning 40.8 68.5 42.0 74.1 35.6 68.1 50.0 76.6 42.1 71.8
Table 7: Performance on the validation set of R2C2-Bench.
Training samples
($\times 4$)
Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
     10,000 38.0 63.7 40.4 71.3 31.2 65.6 44.0 74.4 38.4 68.7
     50,000 40.0 67.2 39.6 71.1 34.0 66.3 49.2 75.6 40.7 70.0
     100,000 40.8 68.5 42.0 74.1 35.6 68.1 50.0 76.6 42.1 71.8
     200,000 39.6 66.9 43.6 73.9 35.4 66.0 49.6 75.6 42.1 70.6
Table 8: Performance of StarCoder-7B with R2C2-Enhanced Tuning on the validation set of R2C2-Bench. Experiments are conducted on R2C2-Bench using different sizes of training data.
5 Conclusion
In this paper, we propose the R2C2-Coder including R2C2-Enhance and R2C2-Bench to enhance and benchmark the real-world repository-level code completion abilities of code LLMs. Specifically, first, R2C2-Enhance is a code prompt construction method, which includes candidate retrieval pool construction and completion prompt construction. Second, the R2C2-Bench is a more challenging repository-level code completion dataset with R2C2-Bench with training, validation, and test splits. Comprehensive results on multiple benchmarks demonstrate the effectiveness of our R2C2-Coder.
6 Acknowledgement
This research project is supported by the Alibaba Aone Copilot Group.
Appendix A Broader Impacts and Limitations
Broader Impacts. In this paper, we propose a methodology with prominent efficiency and ingenious utilization of cross-file information for real-world code completion. We anticipate that this methodology, alongside our fine-tuning strategy and the corresponding benchmarks, will assist developers in enhancing coding efficiency and liberating productivity. Therefore, we hope our work can enable programmers to focus on scenarios with greater creativity.
Limitations. Firstly, to attain enhanced performance, it is essential to fine-tune several hyperparameters. Secondly, despite the need for only 400,000 training samples to achieve outstanding performance, the process demands a certain amount of GPU resources for fine-tuning. Thirdly, four programming languages are covered in our setting, and a large number of languages are not considered in our paper. In our future work, we will continue to investigate to support more languages for facilitating research in this field.
Appendix B More Details
B.1 Details on Abstract Context Extraction
In Fig. 11, we have provided a relatively complex abstract syntax tree for the code in Fig. 10. Besides, we also provide two examples of the extracted abstract context for Python and Java in Fig. 9.
Figure 8: A case found in CrossCodeEval (Ding et al., 2023). As Java supports the package statement, developers could call cross-file APIs in the same package. Such kinds of implicit dependencies are ignored by Ding et al. (2023).
Figure 9: Examples of the generated abstract contexts for Python and Java.
Figure 10: A code example.
Figure 11: An example of the abstract syntax tree generated by the Tree-sitter tool.
B.2 Quality Control Procedure
We begin by establishing the ground-truth completion result as the “reference” and create a set of post-processing filters (i.e., rule-based and model-based rules) to enhance dataset quality. Initially, we eliminate examples based on two criteria: (1) references that are too short (fewer than 10 tokens) and (2) those that are overly long (exceeding 5 lines). Next, we ensure that more than 30% of our references are in multiple lines. To prevent predictable inference from the current file—potentially due to prominent indicators in function names and comments—we input the examples (input prompts) into the DeepSeekCoder-1.3B model (Guo et al., 2024a) for completion and discard any exact matches. This step results in the removal of 57.1%, 39.8%, 48.0%, and 47.2% of samples for Java, Python, TypeScript, and C#, respectively. Lastly, we conducted human annotations on a subset of R2C2-Bench and observed that the dataset possesses sufficient quality for cross-file code completion purposes.
B.3 Dataset Statistics of the validation split of R2C2-Bench
In Table 9, we provide the statistics of the validation split of R2C2-Bench.
Feature Python Java TypeScript C#
# Repositories 23 43 42 21
# Files 250 250 250 250
# Examples 425 717 687 521
Avg. # cross files in prompt 6.03 7.76 6.84 11.28
Avg. # lines in prompt 326.04 333.98 367.46 374.81
Avg. # tokens in prompt 3527.72 3467.47 3451.46 3460.65
Avg. # lines in reference 1.65 1.80 2.12 1.77
Avg. # tokens in reference 16.19 14.64 15.06 13.52
Table 9: The statistics of the validation split of R2C2-Bench.
B.4 Dataset Statistics of CrossCodeEval
As shown in Table 10, we have provided the dataset statistics of the CrossCodeEval benchmark.
Feature Python Java TypeScript C#
# Repositories 471 239 193 99
# Files 1368 745 779 642
# Examples 2665 2139 3356 1768
Avg. # cross files in prompt 3.19 3.15 3.16 3.98
Avg. # lines in prompt 90.6 106.7 116.5 71.1
Avg. # tokens in prompt 938.9 995.3 944.9 584.1
Avg. # lines in reference 1.0 1.1 1.7 1.7
Avg. # tokens in reference 13.2 14.5 17.4 12.5
Table 10: The statistics of the CrossCodeEval.
Appendix C More Experiments
C.1 Details of the baseline models
StarCoder (Li et al., 2023) is a generative decoder-only LM series whose size varies from 1B to 15.5B. Trained on the Stack (Kocetkov et al., 2022) dataset, StarCoder supports a maximum context length of 8K.
DeepSeekCoder (Guo et al., 2024a) is a suite of code models with sizes ranging from 1.3B to 33B trained on a 2-trillion-token manual collected corpus. Benefit from FIM strategy (Bavarian et al., 2022) and Rotary Position Embedding (RoPE) (Su et al., 2024), these models are capable of efficient code generation and infilling under a 16K context window.
Code Llama (Roziere et al., 2023) is a family of large language models based on Llama 2 (Touvron et al., 2023) with 7B, 13B, 34B, and 70B parameters. These models are trained on sequences of 16K tokens and capable of inference on 100K. We conduct experiments with the foundational Code Llama models, not the Python-specialized versions or instruction-tuned ones.
C.2 Details of CrossCodeEval and CrossCodeEval+
CrossCodeEval (Ding et al., 2023) is an arising benchmark for cross-file code completion that is built on a diverse set of real-world, open-sourced, permissively-licensed Github repositories in four popular programming languages: Python, Java, TypeScript, and C#. To create accurate and high-quality test samples, the authors adopt static analysis and filter out those samples solved by starcoderbase-1B within in-file only context.
CrossCodeEval+. We collect the original code bases of CrossCodeEval (Ding et al., 2023) then parse and extract the cross-file dependencies with our R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in CrossCodeEval. Here, we name the reconstructed CrossCodeEval as CrossCodeEval+.
(a) StarCoder
(b) CodeLlama
(c) DeepSeekCoder
Figure 12: Performance on CrossCodeEval after R2C2-Enhanced tuning with various input lengths.
C.3 Scalability of Model Performance on Various Input Lengths
As depicted in Figure 12, we report results produced by code models on CrossCodeEval (Ding et al., 2023) when the input lengths of R2C2-Enhanced tuning range in {512, 1024, 2048, 4096} tokens. There exists a scaling law that the performance is improved when input length increases.
C.4 Execution Time Benchmarking
Execution time measurements are conducted on an Intel Xeon CPU E5-2682 v4 with a base frequency of 2.50 GHz and a turbo boost of up to 4.00 GHz.